OpenStudy (anonymous):

Use elimination to solve the system x - y = 3 and 3x - 8y = 9. What is the value of y?

jhonyy9 (jhonyy9):

x=3+y 3(3+y)-8y=9 9+3y-8y=9 0=5y y=0 x=3